OpenStudy (anonymous):

What is the sum of an 8-term geometric series if the first term is -11, the last term is 859,375, and the common ratio is -5

OpenStudy (texaschic101):

sn = a1(1 - r^n) / (1 - r) n = 8 a1 = -11 r = -5 now we sub s8 = -11(1 - (-5^8) / 1 - (-5) s8 = -11 ( 1 - 390625) / (1 + 5) s8 = -11( - 390624) / 6 s8 = 4296864/6 s8 = 716144